# Empowering Agentic Video Analytics Systems with Video Language Models

### v3

@@ 2.3. Retrieval Augmented Generation
While the videos to be analyzed may span extensive durations, the frames necessary to respond to a specific query are often limited. To validate this observation, we conducted an experiment on VideoMME (Ma et al., 2024) using Qwen2-VL (Wang et al., 2024a). Specifically, we first identified all questions for which Qwen2-VL produced correct answers by uniformly sampling frames from the videos at a rate of 1 FPS. For these questions, we then determined the minimal set of frames required for the VLM to generate the correct answer by iteratively reducing the number of input frames using a binary search strategy  [^1] For example, we initially uniformly sample 100 frames from the video. If the VLM can generate the correct answer based on these frames, we then attempt to reduce the frame set to 50 frames. If 50 frames are still sufficient to produce the correct answer, we further reduce the set to 25 frames. Conversely, if 50 frames are insufficient, we increase the set to 75 frames, iteratively refining the frame selection using binary search strategy.. The results reveal that the frames necessary to answer a given question constitute only a small fraction of the total frames in the video, as shown in Table 1.
@@
Based on this observation, an intuitive approach would be to first retrieve the relevant frames corresponding to a specific query and then generate the final answer based on these frames, a method commonly referred to as Retrieval-Augmented Generation (RAG). To retrieve potentially relevant frames, a straightforward strategy involves vectorized retrieval, where each frame of the video is embedded using a vision-language model, such as CLIP (Radford et al., 2021). At query time, the embedding of the query is used to retrieve relevant frames by comparing the similarity between the query embedding and the vectorized frames.
@@
However, the vectorized retrieval method’s limitations stem from the detailed information contained in the query. Notably, such an approach struggles to handle query-focused summaries (Edge et al., 2024) (e.g., ”What happened in the last few hours?”) or multi-hop queries (Zhao et al., 2024) (e.g., ”What did the man do after he opened the fridge?”), as the retrieved frames often fail to capture key contexts that are not explicitly mentioned in the query descriptions.
@@
To enable effective retrieval, recent research has explored two prominent approaches: video structuring and iterative retrieval. For example, Video-RAG (Luo et al., 2024) structures videos by utilizing various tools to extract information such as audio transcripts (ASR), optical character recognition (OCR) results, and object detection outputs. It then applies RAG techniques to the structured information. However, this method is inherently constrained by the tools employed for video structuring. It is often impractical to predict in advance what types of information need to be extracted and what corresponding tools should be utilized, limiting its adaptability to diverse and dynamic video analytics scenarios.
@@
Alternatively, researchers have proposed obtaining relevant frames through multiple iterative retrieval processes (Wang et al., 2024d, c; Ma et al., 2024; Yang et al., 2024). For instance, VideoAgent (Wang et al., 2024d) typically begins with a coarse-grained sampling of video segments to establish an initial high-level understanding. Based on this, the VLM is prompted to decide which finer-grained segments to retrieve and analyze in subsequent iterations. However, these approaches face significant challenges when applied to video analytics scenarios involving extremely long videos. On one hand, the initial coarse-grained sampling may become insufficient as video length increases, potentially missing critical information. On the other hand, the iterative retrieval and analysis process becomes increasingly computationally expensive as video duration grows, making it impractical for large-scale video analytics tasks.
@@
Recent studies have advanced RAG techniques (Edge et al., 2024; Research, 2024; Guo et al., 2024; Fan et al., 2025) by incorporating knowledge graph construction to enhance the retrieval process. However, these works primarily focus on text-only RAG problems, and adapting such approaches to video analytics remains a significant challenge due to the complexity and multimodal nature of video data. In this paper, we propose Avas, which, to the best of our knowledge, is the first system to enable VLM-powered video analytics by effectively addressing the aforementioned challenges.
3. Avas System Overview
@@
The key idea of Avas lies in leveraging a small VLM to efficiently structure video streams into discrete events, then linking these events by extracting insights from each to construct a comprehensive index. Given a specific query, Avas leverages this index to proactively retrieve relevant information from both the index and the associated raw frames. Ultimately, the retrieved data are utilized by the VLM to produce a coherent and contextually appropriate response.
@@
To build Avas as the L4 video analytics system, we established the following design principles: 1) The analytics should be scalable to any volume of video data, i.e., exceeding hundreds of hours, while ensuring that the computational overhead remains independent of the video length; 2) The index construction must operate in near-real-time, allowing the system to support timely event analytics; 3) The system should accommodate not only fact-based retrieval queries but also query-focused summarization and multi-hop queries, supporting open-ended analytics.
@@
To this end, as depicted in Fig. 1, Avas system is composed of two primary components: near-real-time index construction and agentic retrieval and generation. Within each component, we introduce a set of techniques designed to effectively realize the established design principles.
@@
In the index construction phase, our objective is to design an effective index while ensuring construction efficiency. To achieve this, we introduce the event knowledge graph (EKG) to structure video streams (§4.1). An EKG is a specialized form of a knowledge graph (KG) designed to represent and organize at the granularity of events and their interconnections. Recognizing that events unfold across varying temporal scales, we propose semantic chunking4.2) to extract meaningful events from video streams. Specifically, video streams are segmented into small, uniform chunks (e.g., 3-second intervals), and a small VLM, such as Qwen2.5-VL-7B, is periodically employed to generate detailed content descriptions for these chunks using carefully crafted prompts. Subsequently, neighboring chunks are merged into larger semantic chunks by identifying semantically equivalent textual descriptions with BertScore (Devlin et al., 2019). For each semantic chunk, the small VLM extracts entities and their relationships. Identical entities across different events are linked to ensure consistency and coherence. Ultimately, Avas facilitates the continuous construction of an EKG for a given video stream, regardless of its length, providing a comprehensive representation of semantic events, entities, and their interrelations in near-real-time on typical edge servers.
@@
In the retrieval and generation process, we aim to leverage the constructed index for efficiently retrieving essential and minimal information, and to robustly generate the final answer based on the retrieved data. To achieve this, we first introduce the concept of tri-view retrieval. Specifically, a given query undergoes simultaneous retrieval across three dimensions: events, entities, and visual embeddings. This approach ensures the acquisition of comprehensive and relevant information pertaining to the query. To further support complex queries in L4 VA systems e.g., query-focused summarization and multi-hop queries, we propose an agentic searching mechanism. In particular, by utilizing the LLM as an agent, Avas proactively explores to retrieve additional information from events linked to those retrieved in earlier steps. Avas explores multiple pathways to gather information and formulates a response to the query based on the collected data. Finally, we introduce the thoughts-consistency strategy, which selects the most coherent and accurate final answer from multiple generated candidates.
4. Near-Real-Time Index Construction
4.1. Event Knowledge Graph
@@
Although existing works, such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), utilize knowledge graphs (KGs) to construct retrieval indices, we argue that EKGs are more suitable for video data. The rationale lies in the fundamental difference between the two: KGs focus on static entities (e.g., people, locations, concepts) and their attribute-based relationships, whereas EKGs prioritize modeling dynamic events and their spatiotemporal evolution.
@@
Figure 2. An example of a constructed event knowledge graph and a knowledge graph from wildlife surveillance scenarios for video analytics.
@@
(1) $\mathcal{G}=(\mathcal{E},\mathcal{U},\mathcal{R}),$
where =$\mathcal{E}=\{e_{i}\}_{i=1}^{|\mathcal{E}|}$ represents the temporally ordered set of events, $\mathcal{U}=\{u_{j}\}_{j=1}^{|\mathcal{U}|}$ denotes the entities extracted from the video within each event, and $\mathcal{R}=\mathcal{R}_{ee}\cup\mathcal{R}_{uu}\cup\mathcal{R}_{ue}$ encompasses three types of relationships: 1) temporal event-event relations $\mathcal{R}_{ee}$, such as before and after, which encode temporal logic constraints; 2) semantic entity-entity relations $\mathcal{R}_{uu}$, akin to the relationships found in conventional KGs; and 3) participation relations $\mathcal{R}_{ue}$, which associate entities with their contextual roles within specific events.
4.2. Semantic Chunking
Figure 3. Merging uniform chunks into semantic chunks guided by the pairwise BERTScore distribution.
@@ 4.2. Semantic Chunking
On one hand, large VLMs, such as Qwen2.5-VL-72B, can achieve high accuracy in event detection and transcription, but their substantial computational overhead makes it difficult to process video streams in near-real-time, particularly on resource-constrained edge servers. On the other hand, small VLMs, such as Qwen2.5-VL-7B, offer reduced latency but suffer from performance degradation as the length of the video increases. Furthermore, both large and small VLMs are limited by their constrained context windows. To handle long video content, a common approach is to partition the content uniformly, a process known as chunking. However, events in videos naturally occur at varying and diverse temporal scales. Inaccurate chunking can disrupt the coherence of individual events, thereby increasing the difficulty for VLMs to accurately detect and transcribe them.
@@
To address this, we propose a semantic chunking approach. The core idea involves processing a video stream in the following steps: First, we perform uniform buffering, e.g., dividing the video into fixed-length chunks of 3 seconds each. Next, a small VLM ,e.g., Qwen2.5-VL-7B, is employed to extract representative event descriptions from these chunks with proper prompts. Based on the generated event descriptions, we utilize a text embedding model, such as BERT (Devlin et al., 2019), to measure the similarity between neighboring events. Adjacent events with high similarity are then merged into a single event. Ultimately, this process enables Avas to partition the entire video into semantically meaningful chunks while simultaneously extracting their corresponding descriptions.
@@
Particularly, an input video $V$ is initially divided into uniform chunks $c_{i}$, and a small VLM is employed to generate textual descriptions $d_{i}$ for each chunk $c_{i}$. Subsequently, the semantic similarity between any two uniform chunks is measured by computing the pairwise BertScore (Zhang et al., 2020) for $(d_{i},d_{j})$. Higher similarity scores suggest that the same event may occur across these chunks, making them candidates for semantic merging. Specifically, we adopt two criteria to determine whether certain uniform chunks can be merged into a single semantic chunk: 1) Within a semantic chunk, the similarity between any two uniform chunks must exceed a predefined threshold (e.g., 0.65 in our implementation); 2) After merging, the similarity between the boundaries of adjacent semantic chunks must fall below a sufficiently low threshold. Figure 3 illustrates the semantic chunking process, where a video initially divided into 18 uniform chunks is successfully merged into 9 semantic chunks. Once merged, the small VLM is further utilized to summarize each semantic chunk.
@@
It is important to highlight, although pairwise BertScore computations are performed multiple times, Avas efficiently schedules these computations in parallel, leveraging the hardware parallelism (§6). Consequently, the semantic chunking process does not become a bottleneck in the near-real-time index construction phase, as detailed in §7.
4.3. Entity Extraction and Linking
@@
In addition to extracting event information from videos, Avas also identifies entities and their relationships, as illustrated in Fig. 1. Similar to the approach in (Edge et al., 2024), we utilize a small VLM to extract entities and their relationships from videos using carefully designed prompts for each event.
@@
The identified entities, however, tend to be highly redundant across events within the EKG. Such redundancy not only increases storage requirements but also hampers retrieval efficiency. Thus, it is necessary to de-duplicate and link these entities. Existing works (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025), which primarily focus on text-only RAG problems, typically rely on exact string matching strategies for entity de-duplication. However, in the context of video analytics and EKG, entities are independently extracted from each event by the VLM, leading to potential inconsistencies in entity descriptions for the same concept across different events, e.g.,, ”raccoon” and ”procyon lotor”.
@@
To address this, Avas employs a text embedding model e.g., JinaCLIP (Koukounas et al., 2024), to encode all extracted entities into vector representations. Using embedding similarities as a metric, we apply a standard K-means clustering algorithm to group entities. This approach ensures that semantically similar entities are de-duplicated and linked by forming unified clusters. Subsequently, the centroid of each cluster is used as the representative for the linked entities.
@@
Ultimately, the constructed EKG is stored in a database comprising five tables: events, entities, event-to-event relationships, entity-to-entity relationships, and entity-to-event relationships. Additionally, the raw video frames are vectorized using JinaCLIP (Koukounas et al., 2024) and linked to their corresponding events, enabling comprehensive retrieval in the following phase.
5. Agentic Retrieval and Generation
@@
In the agentic retrieval and generation stage, our primary objectives are to effectively retrieve relevant information using the constructed EKG and to generate robust, contextually accurate responses based on the retrieved data. To achieve this, we introduce an agentic searching mechanism that explores multiple retrieval pathways within the index to gather the necessary information. Additionally, we propose a thought consistency strategy, enabling Avas to generate and evaluate multiple responses from different retrieval pathways, ultimately selecting the most appropriate one.
5.1. Tri-View Retrieval

### v5

@@ 2.3 Retrieval Augmented Generation
While the videos to be analyzed may span extensive durations, the frames necessary to respond to a specific query are often limited. To validate this observation, we conducted an experiment on VideoMME [29] using Qwen2-VL [42]. Specifically, we first identified all questions for which Qwen2-VL produced correct answers by uniformly sampling frames from the videos at a rate of 1 FPS. For these questions, we then determined the minimal set of frames required for the VLM to generate the correct answer by iteratively reducing the number of input frames using a binary search strategy.  [^1] For example, we initially uniformly sample 100 frames from the video. If the VLM can generate the correct answer based on these frames, we then attempt to reduce the frame set to 50 frames. If 50 frames are still sufficient to produce the correct answer, we further reduce the set to 25 frames. Conversely, if 50 frames are insufficient, we increase the set to 75 frames, iteratively refining the frame selection using binary search strategy. The results indicate that, owing to significant temporal redundancy in video data, the frames required to answer a specific query represent only a small fraction of the total frames, as shown in Table 1.
@@
Based on this observation, an intuitive approach would be to first retrieve the relevant frames corresponding to a specific query and then generate the final answer based on these frames, a method commonly referred to as Retrieval-Augmented Generation (RAG). To retrieve potentially relevant frames, a straightforward strategy involves vectorized retrieval, where each frame of the video is embedded using a vision-language model, such as CLIP [33]. At query time, the embedding of the query is used to retrieve relevant frames by comparing the similarity between the query embedding and the vectorized frames.
@@
However, the vectorized retrieval method’s limitations stem from the detailed information contained in the query. Notably, such an approach struggles to handle query-focused summaries [12] (e.g., "What happened in the last few hours?") or multi-hop queries [54] (e.g., "What did the man do after he opened the fridge?"), as the retrieved frames often fail to capture key contexts that are not explicitly mentioned in the query descriptions.
@@
To enable effective retrieval, recent research has explored two prominent approaches: video structuring and iterative retrieval. For example, Video-RAG [28] structures videos by utilizing various tools to extract information such as automatic speech recognition (ASR), optical character recognition (OCR) results, and object detection outputs. It then applies RAG techniques to the structured information. However, this method is inherently constrained by the tools employed for video structuring. It is often impractical to predict in advance what types of information need to be extracted and what corresponding tools should be utilized, limiting its adaptability to diverse and dynamic video analytics scenarios.
@@
Alternatively, researchers have proposed obtaining relevant frames through multiple iterative retrieval processes [45, 47, 29, 48]. For instance, VideoAgent [45] typically begins with a coarse-grained sampling of video segments to establish an initial high-level understanding. Based on this, the VLM is prompted to decide which finer-grained segments to retrieve and analyze in subsequent iterations. However, these approaches face significant challenges when applied to video analytics scenarios involving extremely long videos. On one hand, the initial coarse-grained sampling may become insufficient as video length increases, potentially missing critical information. On the other hand, the iterative retrieval and analysis process becomes increasingly computationally expensive as video duration grows, making it impractical for large-scale video analytics tasks.
@@
Recent studies have advanced RAG techniques [12, 34, 16, 13] by incorporating knowledge graph construction to enhance the retrieval process. However, these works primarily focus on text-only RAG problems, and adapting such approaches to video analytics remains a significant challenge due to the complexity and multimodal nature of video data. In this paper, we propose Ava, which, to the best of our knowledge, is the first system to enable VLM-powered video analytics by effectively addressing the aforementioned challenges.
3 Ava System Overview
@@
The key idea of Ava lies in leveraging a small VLM to efficiently structure video streams into discrete events, then linking these events by extracting insights from each to construct a comprehensive index. Given a specific query, Ava leverages this index to proactively retrieve relevant information from both the index and the associated raw frames. Ultimately, the retrieved data are utilized by the VLM to produce a coherent and contextually appropriate response.
@@
To build Ava as the L4 video analytics system, we established the following design principles: 1) The analytics should be scalable to any volume of video data, i.e., exceeding hundreds of hours, while ensuring that the computational overhead remains independent of the video length; 2) The index construction must operate in near-real-time, allowing the system to support timely event analytics; 3) The system should accommodate not only fact-based retrieval queries but also query-focused summarization and multi-hop queries, supporting open-ended analytics.
@@
To this end, as depicted in Fig. 2, Ava system is composed of two primary components: near-real-time index construction and agentic retrieval and generation. Within each component, we introduce a set of techniques designed to effectively realize the established design principles.
@@
In the index construction phase, our objective is to design an effective index while ensuring construction efficiency. To achieve this, we introduce the event knowledge graph (EKG) to structure video streams (§4.1). An EKG is a specialized form of a knowledge graph (KG) designed to represent and organize at the granularity of events and their interconnections. Recognizing that events unfold across varying temporal scales, we propose semantic chunking4.2) to extract meaningful events from video streams. Specifically, video streams are segmented into small, uniform chunks (e.g., 3-second intervals), and a small VLM, such as Qwen2.5-VL-7B, is periodically employed to generate detailed content descriptions for these chunks using carefully crafted prompts. Subsequently, neighboring chunks are merged into larger semantic chunks by identifying semantically equivalent textual descriptions with BERTScore [11]. For each semantic chunk, the small VLM extracts entities and their relationships. Identical entities across different events are linked to ensure consistency and coherence. Ultimately, Ava facilitates the continuous construction of an EKG for a given video stream, regardless of its length, providing a comprehensive representation of semantic events, entities, and their interrelations in near-real-time on typical edge servers.
@@
In the retrieval and generation process, we aim to leverage the constructed index for efficiently retrieving essential and minimal information, and to robustly generate the final answer based on the retrieved data. To achieve this, we first introduce the concept of tri-view retrieval. Specifically, a given query undergoes simultaneous retrieval across three dimensions: events, entities, and visual embeddings. This approach ensures the acquisition of comprehensive and relevant information pertaining to the query. To further support complex queries in L4 video analytics systems e.g., query-focused summarization and multi-hop queries, we propose an agentic searching mechanism. In particular, by utilizing the LLM as an agent, Ava proactively explores to retrieve additional information from events linked to those retrieved in earlier steps. Ava explores multiple pathways to gather information and formulates a response to the query based on the collected data. Finally, we introduce the thoughts-consistency strategy, which selects the most coherent and accurate final answer from multiple generated candidates.
4 Near-Real-Time Index Construction
4.1 Event Knowledge Graph
@@
Although existing works, such as GraphRAG [12] and LightRAG [16], utilize knowledge graphs (KGs) to construct retrieval indices, we argue that EKGs are more suitable for video data. The rationale lies in the fundamental difference between the two: KGs focus on static entities (e.g., people, locations, concepts) and their attribute-based relationships, whereas EKGs prioritize modeling dynamic events and their spatiotemporal evolution.
@@
Figure 3: An example of a constructed event knowledge graph and a knowledge graph from wildlife monitoring scenarios for video analytics.
@@
$\mathcal{G}=(\mathcal{E},\mathcal{U},\mathcal{R}),$ (1)
where $\mathcal{E}=\{e_{i}\}_{i=1}^{|\mathcal{E}|}$ represents the temporally ordered set of events, $\mathcal{U}=\{u_{j}\}_{j=1}^{|\mathcal{U}|}$ denotes the entities extracted from the video within each event, and $\mathcal{R}=\mathcal{R}_{ee}\cup\mathcal{R}_{uu}\cup\mathcal{R}_{ue}$ encompasses three types of relationships: 1) temporal event-event relations $\mathcal{R}_{ee}$, such as before and after, which encode temporal logic constraints; 2) semantic entity-entity relations $\mathcal{R}_{uu}$, akin to the relationships found in conventional KGs; and 3) participation relations $\mathcal{R}_{ue}$, which associate entities with their contextual roles within specific events.
4.2 Semantic Chunking
@@
On one hand, large VLMs, such as Qwen2.5-VL-72B, can achieve high accuracy in event detection and transcription, but their substantial computational overhead makes it difficult to process video streams in near-real-time, particularly on resource-constrained edge servers [55]. On the other hand, small VLMs, such as Qwen2.5-VL-7B, offer reduced latency but suffer from performance degradation as the length of the video increases. Furthermore, both large and small VLMs are limited by their constrained context windows. To handle long video content, a common approach is to partition the content uniformly, a process known as chunking. However, events in videos naturally occur at varying and diverse temporal scales. Inaccurate chunking can disrupt the coherence of individual events, thereby increasing the difficulty for VLMs to accurately detect and transcribe them.
@@
To address this, we propose a semantic chunking approach. The core idea involves processing a video stream in the following steps: First, we perform uniform buffering, e.g., dividing the video into fixed-length chunks of 3 seconds each. Next, a small VLM, e.g., Qwen2.5-VL-7B, is employed to extract representative event descriptions from these chunks with proper prompts. Based on the generated event descriptions, we utilize a text embedding model, such as BERT [11], to measure the similarity between neighboring events. Adjacent events with high similarity are then merged into a single event. Ultimately, this process enables Ava to partition the entire video into semantically meaningful chunks while simultaneously extracting their corresponding descriptions.
@@
Particularly, an input video $V$ is initially divided into uniform chunks $c_{i}$, and a small VLM is employed to generate textual descriptions $d_{i}$ for each chunk $c_{i}$. Subsequently, the semantic similarity between any two uniform chunks is measured by computing the pairwise BERTScore [50] for $(d_{i},d_{j})$. Higher similarity scores suggest that the same event may occur across these chunks, making them candidates for semantic merging. Specifically, we adopt two criteria to determine whether certain uniform chunks can be merged into a single semantic chunk: 1) Within a semantic chunk, the similarity between any two uniform chunks must exceed a predefined threshold (e.g., 0.65 in our implementation); 2) After merging, the similarity between the boundaries of adjacent semantic chunks must fall below a sufficiently low threshold. Fig. 4 illustrates the semantic chunking process, where a video initially divided into 18 uniform chunks is successfully merged into 9 semantic chunks. Once merged, the small VLM is further utilized to summarize each semantic chunk.
@@
It is important to highlight, although pairwise BERTScore computations are performed multiple times, Ava efficiently schedules these computations in parallel, leveraging the hardware parallelism (§6). Consequently, the semantic chunking process does not become a bottleneck in the near-real-time index construction phase, as detailed in §7.
Figure 4: Merging uniform chunks into semantic chunks guided by the pairwise BERTScore distribution.
4.3 Entity Extraction and Linking
@@
In addition to extracting event information from videos, Ava also identifies entities and their relationships, as illustrated in Fig. 2. Similar to the approach in [12], we utilize a small VLM to extract entities and their relationships from videos using carefully designed prompts for each event.
@@
The identified entities, however, tend to be highly redundant across events within the EKG. Such redundancy not only increases storage requirements but also hampers retrieval efficiency. Thus, it is necessary to de-duplicate and link these entities. Existing works [12, 16, 13], which primarily focus on text-only RAG problems, typically rely on exact string matching strategies for entity de-duplication. However, in the context of video analytics and EKG, entities are independently extracted from each event by the VLM, leading to potential inconsistencies in entity descriptions for the same concept across different events, e.g.,, "raccoon" and "procyon lotor".
@@
To address this, Ava employs a text embedding model e.g., JinaCLIP [22], to encode all extracted entities into vector representations. Using embedding similarities as a metric, we apply a standard K-means clustering algorithm to group entities. This approach ensures that semantically similar entities are de-duplicated and linked by forming unified clusters. To represent each cluster, we compute the centroid of the embedding vectors of all entities within that cluster, which serves as the representative feature of the merged entity.
@@
Ultimately, the constructed EKG is stored in a database comprising five tables: events, entities, event-to-event relationships, entity-to-entity relationships, and entity-to-event relationships. Additionally, the raw video frames are vectorized using JinaCLIP [22] and linked to their corresponding events, enabling comprehensive retrieval in the following phase.
5 Agentic Retrieval and Generation
@@
In the agentic retrieval and generation stage, our primary objectives are to effectively retrieve relevant information using the constructed EKG and to generate robust, contextually accurate responses based on the retrieved data. To achieve this, we introduce an agentic searching mechanism that explores multiple retrieval pathways within the index to gather the necessary information. Additionally, we propose a thought consistency strategy, enabling Ava to generate and evaluate multiple responses from different retrieval pathways, ultimately selecting the most appropriate one.
5.1 Tri-View Retrieval